OpenStudy (preetha):

How do you solve this equation: 3x2+4x-7=12 that is 3x^2

OpenStudy (misty1212):

HI!!`

OpenStudy (preetha):

Hi

OpenStudy (misty1212):

start by subtracting \(12\) from both sides to get \[x^2+4x-19=0\] then use the quadratic formula since this one is not going to factor for you

OpenStudy (michele_laino):

we have to apply the standard formula, namely: \[3{x^2} + 4x - 19 = 0\]

OpenStudy (misty1212):

oops i meant \[3x^2+4x-19=0\]

OpenStudy (misty1212):

maybe it does factor, who knows?

OpenStudy (misty1212):

nope, quadratic formula is what you need do you know it?

OpenStudy (preetha):

Hmmm. Keep going. I am trying to keep up. (-:

OpenStudy (michele_laino):

then: \[\large x = \frac{{ - 2 \pm \sqrt {4 + 57} }}{3} = \frac{{ - 2 \pm \sqrt {61} }}{2}\]

OpenStudy (alexandervonhumboldt2):

lol i think preetha is just testing the system and checking how QH's work.

OpenStudy (misty1212):

solution to \[ax^2+bx+c=0\] is \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (preetha):

What do I use for b and a etc?

OpenStudy (misty1212):

in your example \[a=3,b=4,c=-19\]

OpenStudy (misty1212):

first step is substitution you get \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\\ x=\frac{-4\pm\sqrt{4^2-4\times 3\times (-19)}}{2\times 3}\]

OpenStudy (misty1212):

it is arithmetic from there on in

OpenStudy (michele_laino):

oops.. I have applied the subsequent formula: \[\large x = \frac{{ - \left( {b/2} \right) \pm \sqrt {{{\left( {b/2} \right)}^2} - ac} }}{a}\]

OpenStudy (preetha):

Thanks Misty. Thanks Michele. You folks use the equation editor so well. I am embarrassed by my equation. Michele, I will go and rate you help now so you can get the OwlBucks reward.

OpenStudy (misty1212):

lol thanks, i need a few more owls

OpenStudy (michele_laino):

Thanks @Preetha Thanks @misty1212

OpenStudy (isaiah.feynman):

Could also complete the square. I like it.

OpenStudy (michele_laino):

oops.. I have made an error, the right answer is: \[\large x = \frac{{ - 2 \pm \sqrt {4 + 57} }}{3} = \frac{{ - 2 \pm \sqrt {61} }}{3}\]

OpenStudy (preetha):

Thanks Michele and Misty. You are awesome helpers.

OpenStudy (michele_laino):

thanks again :) @Preetha

OpenStudy (nincompoop):

That OR show some algebra skills by completing the square :)

OpenStudy (preetha):

Nincompoop, thanks.

OpenStudy (isaiah.feynman):

See nincompoop agrees with me. :P